Circular No. 7871

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cc IN ANONYMOUS GALAXY W. D. Li, University of California at Berkeley, reports the LOTOSS (cf.

IAUC 7514) discovery of an apparent supernova (mag ∼ 18.5) on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Apr. 4.2 and 9.3 UT. SN 2002cc is located at α = 8^{h}34^{m}01^{s}.97, δ = +55^{o}39′12′′.4 (equinox 2000.0), which is 1′′.2 west and 5′′.3 north of the nucleus of an anonymous galaxy located ∼ 65′′ west and 73′′ north of UGC 4470. A KAIT image of the same field on Mar. 25.3 UT showed nothing at this position (limiting mag ∼ 19.5).

T. Boles, Coddenham, Suffolk, reports his independent discovery of this supernova at mag 17.2 on a 30-s unfiltered CCD image taken on Apr. 8.876 UT (limiting mag 18.5) with a 0.35-m Schmidt-Cassegrain reflector, in the course of the U.K. Nova/Supernova patrol. He provides end figures 02^{s}.11, 12′′.5 and notes that the object is not present on his images from 2001 Dec. 10 (limiting mag 19.0), 2002 Jan. 17 or Feb. 16. It is also not present on Palomar Sky Survey images taken on 1997.908 (red, limiting mag 21.0) and 1991.110 (blue, limiting mag 20.5).

SUPERNOVA 2002cd IN NGC 6916 M. Armstrong, Rolvenden, Kent, reports his discovery of an apparent

supernova (mag 17.5) on an unfiltered CCD image (limiting mag 19.0) taken with a 0.35-m Schmidt-Cassegrain telescope on Apr. 8.157 UT in the course of the U.K. Nova/Supernova patrol. The new object is confirmed at mag 17.0 on an image taken on Apr. 9.04 and is located at α = 20^{h}23^{m}34^{s}.42, δ = +58^{o}20′47′′.4 (equinox 2000.0), or ∼ 10′′ east and 10′′ north of the center of NGC 6916. The new object is not present on Armstrong’s images from 2001 July 15 (limiting mag 19.5), Nov. 1 and Nov. 10 (both limiting mag 19.0), or on Palomar Sky Survey images taken on 1989.573 (blue, limiting mag 22.5) and 1991.527 (red, limiting mag 20.8).

SUPERNOVA 2002bt IN UGC 8584 T. Matheson, S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002bt (cf. IAUC 7862), obtained by M. Calkins on Apr. 3.42 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova two weeks after maximum.

2002 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden